THE DEPENDENCE OF INTERMOLECULAR INTERACTIONS UPON VALENCE COORDINATE EXCI-TATION: THE v_{HF} =4 LEVELS OF ArHF

\underline{CHENG-CHI CHUANG}, and WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, MA 02138.

The valence state dependence of the Ar−HF interaction potential is extended to v_{HF} = 4. Three new ArHF (v_{HF} = 4) states, (4000), (4100), and (4110), are observed between 14780 and 14880 cm^{−1} using intracavity laser induced fluorescence. The spectroscopic constants of these states are: (4000) ν_{0} = 14783.60323(30) cm^{−1}, and B = 0.1036068(68) cm^{−1}; (4100) ν_{0} = 14867.41906(70) cm^{−1}, B = 0.102612(27) cm^{−1}; and (4110) ν_{0} =14875.04673(31) cm^{−1}, B = 0.1012823(73) cm^{−1} respectively. In conjunction with the previous ArHF results (v_{HF} =0−3), the spectral red shifts of ArHF(v000) increase more rapidly than linearly, from 9.654 cm^{−1} at v=1 to 48.024 cm^{−1} at v=4. The rotational constants of ArHF(v000) increase essentially linearly with v_{HF}, noticeably increased by 1.30 % (40 MHz) at v=4. The classical turning point of HF is extended by 0.4 Å from r_{e} at v=4, showing no evidence for Ar−H repulsion. The spectral red shift for linear hydrogen bonded Ar−HF (v000) indicates a strong enhancement of binding energy upon HF valence bond elongation, while the rotational constant indicate an almost surprising decrease in heavy atom separation. Both the T-shaped ArHF(v110) and anti-linear Ar−FH(v100), however, show very little dependence of binding energy upon v_{HF} valence excitation. These observations are in good accord with the ab initio calculations of the Ar−HF intermolecular potential surface.